A heat engine does 30.0 J of work per cycle while
Answer #1

Work done by the heat engine =W=30J

Waste heat =Q_{c}=40J

Heat input =Q_{H}

Q_{H}=Q_{c}+W

Q_{H}=(40+30)J=70J

Efficiency =n=\frac{W}{Q_{H}}=\frac{30}{70}=0.4286

Hence the efficiency is 42.86 %